Heres the problem 4(5n-7)=10n+2 Ive tried it a couple of times but i always get stuck here
4(5n-7)=10n+2
20n-28=10n+2
30n-26-------This is where i get stuck

Answer by Born2TeachMath(20)

You're good through line #3:
20n - 28 = 10n + 2.
At this point, the idea is to separate the terms with variables on one side and the numbers on the other. So, subtract 10n from each side, and add 28 to each side.
20n - 28 + 28 - 10n = 10n - 10n + 2 + 28
The -28 and +28 cancel out on the left side, and the 10n and -10n cancel out on the right side.
20n - 10n = 2 + 28
Simplifying what's left leaves:
10n = 30.
Then divided each side by 10 to get rid of the 10 on the left side.
10n/10 = 30/10.
n = 3 is your answer!
REMEMBER, EQUATIONS HAVE TWO SIDES! YOU ALWAYS WILL HAVE TWO SIDES. USE THEM TO SEPARATE THE VARIABLES FROM THE NUMBERS, AND THEN DIVIDE OUT THE LAST COEFFICIENT TO SOLVE FOR THE VARIABLE.
This method works for linear equations only (equations with x^1 powers only). If there are other powers, then other methods must be used...